### v1

# Greenberger-Horne-Zeilinger state is the best probe for multiparameter estimation of independent local fields

Abstract
We show that estimation of multiple independent field strengths of a local Hamiltonian, i.e., one formed by a sum of single-party terms, necessitates the utility of a highly symmetric genuine multiparty entangled state, viz. the Greenberger-Horne-Zeilinger (GHZ) state, as the input probe. This feature however depends on the choice of the weight matrix considered, to define a figure of merit in the multiparameter estimation. We obtain this result by providing a lower bound on the precision of multiparameter estimation, optimized over input probes, for an arbitrary weight matrix. We find that the bound can be expressed in terms of eigenvalues of the arbitrary multiparty local encoding Hamiltonian. We show that there exists a weight matrix for which this bound is attainable only when the probe is the GHZ state, up to a relative phase. In particular, no pure product state can achieve this lower bound. Indeed, the gap - in precision - between genuinely multiparty entangled and product states acting as probes, increases with increasing number of parties. Finally we also prove that using a probe that is in a mixed state provides a precision lower than that for the GHZ state. To emphasize the importance of the weight matrix considered, we also prove that the choice of identity operator as the same - thereby ignoring the covariances in the precision matrix - does not require the use of genuine multiparty entanglement in input probes for attaining the best precision, and the optimal probe can be a pure product.
I introduction
The goal of quantum metrology is to estimate one or more parameters that are encoded via some physical operation onto a system, and to demonstrate an enhancement in the estimation, fueled by one or more quantum resources. Enhancement in precision of estimated parameters due to the presence of quantum resources is a fundamental tenet of quantum-enhanced sensing, and has been widely analyzed both theoretically and experimentally Giovannetti et al. (2011); Tóth and Apellaniz (2014); Demkowicz-Dobrzański et al. (2015); Tan and Jeong (2019). Most of these analyses involve the estimation of a single parameter, while recent advancements have been made on the quantum-enhanced estimation of multiple parameters simultaneously Vidrighin et al. (2014); Ragy et al. (2016a); Pezzè et al. (2017); Yousefjani et al. (2017); Gessner et al. (2018a); Zhuang et al. (2018); Gessner et al. (2018b); Yang et al. (2019); Liu et al. (2019); Goldberg et al. (2020); Gross and Caves (2020a, b); Kull et al. (2020); Lu and Wang (2021); Goldberg et al. (2021); Yang et al. (2022); Górecki and Demkowicz-Dobrzański (2022); Miyazaki and Matsumoto (2022); Reilly et al. (2023). Experimental progress in this direction can e.g. be found in Spagnolo et al. (2012). Metrological advantage induced by quantum resources can be incorporated into the system, for example, though the utilization of entanglement in input probes or measurement operators, or both. It is useful to look into the minimal standard deviation or “uncertainty” of the estimator in assessment of a single parameter to find the best metrological precision. There is a fundamental bound on the standard deviations of estimators, given in terms of a quantity referred to as the Fisher information Wootters (1981); Braunstein and Caves (1994). The Fisher information, in a way, indicates the amount of information about the unknown parameter that can be decoded through the process of estimation. If the Fisher information is maximized over all possible measurements, it gives the quantum Fisher information, which leads to the quantum Cramér-Rao bound.
Here we analyze the situation where multiple (local) field strengths of a many-body encoding Hamiltonian are simultaneously estimated, with each field strength being independent and, in principle, distinct from the others. This is therefore a scenario of multiparameter estimation, and unlike the case of assessing a single parameter where variance is a useful quantity, the covariance matrix of the estimators is here the object of interest. The simultaneous estimation of independent local fields of a many-body encoding Hamiltonian is an arena which is hitherto unexplored. We find that even though the parameters to be estimated are independent local ones, only the Greenberger-Horne-Zeilinger (GHZ) M. Kafatos (Ed.); Mermin state, up to a relative phase, in the eigenbasis of the encoding Hamiltonian, is the optimal probe. To measure the precision of estimation of multiple parameters, and to compare with other instances, it is usually convenient to consider a trace of the matrix inequalities after taking a product with a positive semi-definite weight matrix. The minimal value of this quantity provides the best metrological precision in multiparameter estimation for the chosen weight matrix Magdalena Szczykulska and Datta (2016); Yang et al. (2019); Hamann et al. (2023). We prove the result about optimality and uniqueness of the GHZ probe in three steps. First, we optimize the quantum Cramér-Rao bound with respect to pure probes, and provide a general lower bound in terms of an arbitrary weight matrix. The bound holds for an arbitrary local encoding Hamiltonian and for an arbitrary number of parties. Then we show that for a particular choice of the weight matrix and for a given local encoding Hamiltonian, the bound is independent of the number of parties. In the second step, we prove that, when optimized over pure input probes, the GHZ state, up to a relative phase, is the only state that attains the bound. Finally, in the third step, we have scanned the space of all states, including mixed ones, and shown that no mixed state provides a precision better than that provided by the GHZ state. Furthermore, we find the best precision attainable with pure product probes, and prove that the precision, in this case, worsens with increasing number of probes. Thereby we show that the gap in the figure of merit between optimal product and GHZ probes gradually increases with increasing number of parties. Along the way of proving our main results, we provide an axiomatic formulation of why one can choose the trace of the product of a semi-definite matrix with the quantum Fisher information matrix as a “figure of merit” in quantum multiparameter estimation.
The problem of estimation of the common field strength of an encoding Hamiltonian has been carefully explored. This is one of the fundamental problems that has been dealt with in single-parameter quantum metrology. For instance, it is known that while estimating the common field strength of an encoding Hamiltonian whose local field terms are Pauli-$z$, the GHZ state turns out to be the best probe - providing better scaling of the minimum error than optimal product ones Giovannetti et al. (2006). Here we prove that even if we estimate multiple independent field strengths of a local Hamiltonian, then also the optimal probe is the GHZ, up to a relative phase. Estimation of the common field strength - common over parties - of an encoding Hamiltonian is widely explored Skotiniotis et al. (2015); Ouyang and Rengaswamy (2023); Deng et al. (2021); Bhattacharyya et al. (2024); Salvia et al. (2023); Su et al. (2024), while the simultaneous estimation of non-commuting common fields is less delved into. For instance, the metrological protocol for multiparameter estimation of field strengths corresponding to non-commuting but common fields is examined both in noiseless Baumgratz and Datta (2016) and noisy scenarios Ho et al. (2020); Le et al. (2023).
The remainder of the paper is arranged as follows. In Sec. II, the preliminaries of quantum multiparameter estimation are discussed. In Sec. III, we provide a general formulation of intuitively satisfactory axioms that are to be fulfilled by a figure of merit used in the estimation of multiple parameters. Sec. IV comprises one of our main results. Here we provide a general lower bound in the estimation of independent local fields of an encoding Hamiltonian. As a special case, we prove that the minimum value of the figure of merit is a constant when the local (single-body) part of the many-body encoding Hamiltonian is Pauli-$z$ and a particular weight matrix is chosen. Followed by this, we prove that in this case, the highest sensitivity is attained only by using the GHZ state, up to a relative phase, as the probe, and all product states provide a lower precision. Interestingly, we find that the gap in precision between GHZ and optimal product probes increases with an increase in the number of parties. In Sec. V, we consider arbitrary input probes, including mixed ones, and prove that the optimal sensitivity is only provided by pure states. Sec. VI deals with the situation of attainment of the lower bound of the figure of merit if the weight matrix is considered to be identity. Finally, the concluding remarks are presented in Sec. VII.
II Multiparameter estimation
Let a set of parameters, $\phi=\{\phi_{1},\phi_{2},...,\phi_{n}\}$, be encoded onto an input probe, $\rho$, using some physical process. The aim is to estimate these parameters with best achievable precision by performing suitable measurements, $M_{x}$, on the encoded probe, $\rho_{\phi}$. The probabilities of the measurement outcomes, $x$, belong to to the distribution, $p(x|\phi)=\text{Tr}(\rho_{\phi}M_{x})$. Based on these measurement outcomes, one can calculate the precision in estimating the parameters by assigning estimator functions corresponding to each parameter, where the set of estimator functions is given by $\hat{\phi}=\{\hat{\phi_{1}},\hat{\phi_{2}},...,\hat{\phi_{n}}\}$. These estimators are all assumed to be unbiased, i.e. $\langle\hat{\phi}_{i}\rangle=\phi_{i}$, $\forall i$, where the average is taken with respect to the probability distribution, $p(x|\phi)$.
The problem of estimation of multiple parameters is formulated in terms of the covariance matrix, whose elements are defined as $\text{Cov}(\hat{\phi})_{ij}=\langle(\hat{\phi}_{i}-\phi_{i})(\hat{\phi}_{j}-\phi_{j})\rangle$, $\forall i,j$ where the expectation is again taken with respect to the probability distribution, $p(x|\phi)$. One of the fundamental bounds in the estimation of a single parameter, referred to as the Cramér-Rao bound, is given in terms of the Fisher information. In the estimation of more than one parameters, the corresponding multiparameter Cramér-Rao bound bound is obtained in terms of an $n\times n$ Fisher information matrix, $F$, whose elements are defined in terms of the probability distribution, $p(x|\phi)$, as
$F_{ij}=\sum_{x}p(x|\phi)\left(\frac{\partial\ln p(x|\phi)}{\partial\phi_{i}}\right)\left(\frac{\partial\ln p(x|\phi)}{\partial\phi_{j}}\right)\;\;\forall i,j.$ (1)
The Fisher information matrix, $F$, is related to the covariance matrix for a set of unbiased estimators, $\hat{\phi}$, by the multiparameter Cramér-Rao bound, given by
$\nu\text{Cov}(\hat{\phi})\geq F^{-1},$ (2)
where $\nu$ is the number of times the experiment is repeated. As the multiparameter Cramér-Rao bound is given in terms of matrices, we need to identify a suitable figure of merit for estimating the metrological precision of a set of parameters to compare it with other scenarios. It is convenient to consider a trace of the matrix inequalities after taking a product with a positive semi-definite weight matrix, $W$. Depending upon the choice of the weight matrix, a linear sum of the uncertainties of and the correlations between the different estimators play the role of a figure of merit. It follows from inequality (2) that for a $W\geq 0$, the quantity, $\nu\text{Tr}(W\text{Cov}(\hat{\phi}))\geq\text{Tr}(WF^{-1})$. Similar to the single-parameter case, the multiparameter Cramér-Rao bound is saturated by the maximum likelihood estimator in the asymptotic limit, i.e. when the number of repetitions of the experiment is very large Kay (1993).
The right hand side of the inequality, $\nu\text{Tr}(W\text{Cov}(\hat{\phi}))\geq\text{Tr}(WF^{-1})$, depends on the choice of probe state and the measurements performed. In principle, one can also minimize the quantity, $\text{Tr}(WF^{-1})$, over all measurements and obtain the quantum Cramér-Rao bound in terms of the quantum Fisher information matrix (QFIM), which we refer to as $F_{Q}$. The corresponding quantum Cramér-Rao inequality is given by
$F^{-1}\geq F_{Q}^{-1}.$ (3)
The elements of QFIM, $F_{Q}$, are defined as
$(F_{Q})_{ij}=\frac{1}{2}\text{Tr}(\rho_{\phi}\{L_{i},L_{j}\}),\;\;\;\forall i,j$ (4)
where the curly brackets denote anti-commutator, and $L_{i}$ is a Hermitian matrix, referred to as the symmetric logarithmic derivative (SLD) corresponding to the encoded state, $\rho_{\phi}$. It is defined using the relation $\frac{1}{2}(L_{i}\rho_{\phi}+\rho_{\phi}L_{i})=\partial_{\phi_{i}}\rho_{\phi}$, for any encoded $\rho_{\phi}$. It is to be noted that the QFIM is real and symmetric, i.e. $(F_{Q})_{ij}=(F_{Q})_{ji}$. The SLD operator corresponding to the $i^{\text{th}}$ parameter is explicitly given by
$L_{i}=2\sum_{a,b}\frac{\bra{\psi_{a}}\partial_{\phi_{i}}\rho_{\phi}\ket{\psi_{b}}}{\lambda_{a}+\lambda_{b}}|\psi_{a}\rangle\langle\psi_{b}|,$ (5)
where the encoded state is written in spectral decomposition as $\rho_{\phi}=\sum_{k}\lambda_{k}|\psi_{k}\rangle\langle\psi_{k}|$. If the encoded state is pure, say |$\ket{\psi_{\phi}}$, then the corresponding SLD operator for the $i^{\text{th}}$ parameter further simplifies to
$L_{i}=2(|\partial_{\phi_{i}}\psi_{\phi}\rangle\langle\psi_{\phi}|+|\psi_{\phi}\rangle\langle\partial_{\phi_{i}}\psi_{\phi}|).$ (6)
The diagonal elements of the QFIM are given by $(F_{Q})_{ii}=\text{Tr}(\rho_{\phi}L_{i}^{2})$, which essentially provides the quantum Fisher information for the $i^{\text{th}}$ parameter.
It is known that the optimal measurement for a given parameter can be constructed from projectors onto the eigenbasis of SLD corresponding to that parameter. However, unlike the Cramér-Rao bound in (2), the quantum Cramér-Rao bound in (3) is not always saturable in case of multiparameter estimation. This is because the projectors onto the eigenspace of $L_{i}$ may not commute with those of $L_{j}$, where $i\neq j$, $\forall i,j$. Therefore it may not be possible to find a single measurement that saturates the the quantum Cramér-Rao bound in (3). It is however to be noted that the condition, $[L_{i},L_{j}]=0$, $\forall i,j$ is only necessary and not sufficient to saturate the quantum Cramér-Rao bound. A more general necessary and sufficient criterion of saturability of multiparameter quantum Cramér-Rao bound has been provided for both pure states Matsumoto (2002) and mixed states Ragy et al. (2016b). This condition is given by
$\text{Tr}(\rho_{\phi}[L_{i},L_{j}])=0\;\;\forall i,j,$ (7)
where not the commutator itself, but its expectation value with respect to the encoded state vanishes.
The actual quantities of interest that are henceforth considered to measure and compare the sensitivities of a set of parameters, $\phi$, satisfy the following inequalities:
$\nu\text{Tr}(W\text{Cov}(\hat{\phi}))\geq\text{Tr}(WF^{-1})\geq\text{Tr}(WF_{Q}^{-1}),$ (8)
where $W$ is an arbitrary positive semi-definite matrix. The first inequality in (8) is obtained by optimizing over all unbiased estimators, while the second inequality is obtained after an optimization over all measurements. The quantity, $\text{Tr}(WF_{Q}^{-1})$, therefore, still depends on the choice of the input probe and the encoding.
III Axiomatic formulation for figure of merit in multiparameter estimation
In this section, we formally consider a figure of merit that is used to gauge the metrological precision in the estimation of a set of unknown parameters. We show that this figure of merit satisfies certain axioms that should be fulfilled by a suitable measure of estimation precision in multiparameter quantum metrology. We consider the quantum Fisher information-based figure of merit to be given by $\text{Tr}(WF_{Q}^{-1})$, where $F_{Q}$ is the QFIM, and $W$ is an arbitrary positive semi-definite weight matrix. The quantity, $\text{Tr}(WF_{Q}^{-1})$, can be further minimized over the choice of input probes. We here provide the axioms that are satisfied by the quantity, $\text{Tr}(WF_{Q}^{-1})$, which certifies its use as a valid figure of merit for multiparameter estimation.
Positivity: QFIM is positive semi-definite, i.e. all its eigenvalues are non-negative. Therefore, provided that the inverse of $F_{Q}$ exists, $F_{Q}^{-1}$ has all its eigenvalues equal the inverse of those of $F_{Q}$, and are therefore all non-negative. This means $F_{Q}^{-1}\geq 0$, and since $W\geq 0$, this implies that the quantity, $\text{Tr}(WF_{Q}^{-1})\geq 0$.
Monotonicity: This axiom is based on our belief that application of noise on the encoded probe will deteriorate the sensing precision. QFIM is monotonic under the action of CPTP operations on the encoded probe Liu et al. (2019), i.e.
$F_{Q}(\rho_{\phi})\geq F_{Q}\left(\Lambda({\rho_{\phi}})\right),$ (9)
where $\Lambda$ is an arbitrary CPTP map acting on the encoded probe, $\rho_{\phi}$. Now let us find whether the quantity, $\text{Tr}(WF_{Q}^{-1}({\rho_{\phi}}))$ is monotonic under CPTP operations. Since the inverse of QFIM is also positive semidefinite, if the inverse matrices exist, we can pre-multiply both sides of the inequality (9) by $F_{Q}(\rho_{\phi})^{-1}$ and post-multiply both sides by $F_{Q}(\Lambda({\rho_{\phi}}))^{-1}$, which gives
$F_{Q}(\Lambda({\rho_{\phi}}))^{-1}\geq F_{Q}(\rho_{\phi})^{-1}.$ (10)
Now for any $W\geq 0$, we obtain the following inequation:
$\text{Tr}\left[WF_{Q}(\rho_{\phi})^{-1}\right]\leq\text{Tr}\left[WF_{Q}(\Lambda({\rho_{\phi}}))^{-1}\right].$ (11)
Therefore we find that the figure of merit given by $\text{Tr}(WF_{Q}^{-1}({\rho_{\phi}}))$, is monotonic under the action of CPTP operations. This indicates in particular that under the action of a noise, which will be a CPTP channel, on the encoded probe, the amount of information extractable from the process, as quantified by the QFIM, gets diminished, thereby deteriorating metrological precision.
IV Estimation of independent local fields using pure probes
We consider unitary encoding of probe states. The encoding unitary operator is given by $U=\exp(-i\kappa\widetilde{H}t/\hbar)$, where $\kappa$ has the dimension of $time^{-1}$. We observe the system at time, $t=1$. The generator of the unitary contains an arbitrary $N$-party many-body encoding Hamiltonian, $\widetilde{H}$, containing single-body terms. Our aim is to simultaneously estimate all the field strengths corresponding to every local term in the Hamiltonian. We provide a general lower bound of the quantum Fisher information-based figure of merit, in terms of $\widetilde{H}$ and an arbitrary weight matrix, $W$, in the initial part of this subsection. In the succeeding subsections IV.1 and IV.2, our results pertain to a special choice of the Hamiltonian, $\widetilde{H}=H$, and weight matrix, $W=\mathcal{W}$. There, every single-body term of the many-body Hamiltonian, $H$, is considered to be the Pauli-$z$ operator. Such an Hamiltonian is given by
$H=\hbar\sum_{i=1}^{N}h_{i}\sigma_{z}^{i},$ (12)
where the notation, $\sigma_{z}^{i}$, denotes that the $\sigma_{z}$ operator acts on the $i^{\text{th}}$ party. The dimensionless quantity, $h_{i}$, is the local field strength corresponding to the $i^{\text{th}}$ party. Every $h_{i}$ is assumed to be independent of the other for all values of $i$. Our aim is to simultaneously estimate these local field strengths, $h_{i}$. Let us consider an input probe, |$\ket{\psi_{0}}$, which evolves unitarily under the action of $U$, and let the encoded state after unitary evolution be |$\ket{\psi_{h}}$. The subscript, $h$, denotes that the encoded state depends on the field strengths, $h=\{h_{1},h_{2},...,h_{N}\}$. Following the encoding, a measurement is performed on the encoded probe. The outcomes of measurement follow certain probabilities, from which the Fisher information matrix is obtained, which satisfies the Cramér-Rao inequality. Optimization of the Fisher information matrix-based figure of merit with respect to all measurements generates the quantum Cramér-Rao bound, given in terms of the QFIM. However, in multiparameter estimation, there may not always exist a single measurement that saturates the quantum bound. In our case, however, the SLD operators corresponding to each $h_{i}$ satisfy Eq. (7). Hence in this case, there exists a common eigenbasis of all the SLD operators, and projectors onto those eigenvectors saturate the quantum Cramér-Rao bound (3). So the quantum Cramér-Rao bound, in this case, is attainable and it is reasonable to consider the quantum Fisher information-based figure of merit. This is given by $\text{Tr}(W\mathcal{F}_{Q}^{-1})$, where $W$ is an arbitrary positive semi-definite weight matrix and $\mathcal{F}_{Q}$ is the QFIM corresponding to the encoding Hamiltonian, $H$. The elements of the $N\times N$ matrix, $\mathcal{F}_{Q}$, are given by
$(\mathcal{F}_{Q})_{ii}$ $=$ $4\left[1-\bra{\psi_{0}}\sigma_{z}^{i}\ket{\psi_{0}}^{2}\right],\;\;\forall i$ (13)
$(\mathcal{F}_{Q})_{ij}$ $=$ $4\Big{[}\bra{\psi_{0}}\sigma_{z}^{i}\otimes\sigma_{z}^{j}\ket{\psi_{0}}$
[…]
$=$ $(\mathcal{F}_{Q})_{ji},$
where each of $i$ and $j$ can take values $1$ to $N$. The elements of the QFIM in equations (13) and (IV) originally involved expectations with respect to the encoded state, |$\ket{\psi_{h}}$. However, since $[U,\sigma_{z}^{i}]=[U,\sigma_{z}^{i}\otimes\sigma_{z}^{j}]=0$, these expectations can equivalently be taken with respect to |$\ket{\psi_{0}}$, i.e. the input probe. The quantity, $(\mathcal{F}_{Q})_{ii}$, is four times the variance of $\sigma_{z}^{i}$, when the variance is taken with respect to the probe state. This is essentially the quantum Fisher information that one would obtain if only the $i^{\text{th}}$ parameter would have been estimated, considering the others to be known. The term, $(\mathcal{F}_{Q})_{ij}$, on the other hand, is the covariance between $\sigma_{z}^{i}$ and $\sigma_{z}^{j}$, where the covariance is taken with respect to the probe state. This effectively gives the “correlation” or “coupling” between the errors associated with $h_{i}$ and $h_{j}$. It is evident from equations (13) and (IV) that the diagonal terms of the QFIM, $(\mathcal{F}_{Q})_{ii}$, can vary between 00 and $1$, whereas, the off-diagonal terms, $(\mathcal{F}_{Q})_{ij}$ can vary from $-1$ to $1$. We consider a multiparameter estimation scheme where the contribution of both the diagonal and off-diagonal terms in QFIM are taken into account, and consider an equal weightage of all the terms in the QFIM. This motivates us to choose a weight matrix $\mathcal{W}=|\mathcal{W}\rangle\langle\mathcal{W}|$, where |$\ket{\mathcal{W}}=\sum_{i=0}^{N-1}\ket{i}$ in the computational basis. This specific weight matrix is positive semi-definite, and satisfies the two axioms of positivity and monotonicity provided in the preceding section.
In theorem 1 of this section, we consider the quantum Fisher information-based figure of merit, given by $\text{Tr}(WF_{Q}^{-1})$, and minimize it with respect to all $N$-party pure input probes for an arbitrary choice of $W$, and an arbitrary encoding many-body Hamiltonian, $\widetilde{H}$, comprising of a sum of single-body terms. Then we show that specifically for $W=\mathcal{W}$ and $F_{Q}=\mathcal{F}_{Q}$, the lower bound is a constant and independent of the value of the number of parties. Additionally, we find the condition of saturability of the lower bound that we provide. Further, in subsection IV.1, still considering the weight matrix to be $\mathcal{W}$ and the QFIM to be $\mathcal{F}_{Q}$, we find that optimal precision is attainable using only the Greenberger-Horne-Zeilinger state, up to a relative phase, when the optimization is performed over all pure states. In subsection IV.2, we find the minimum value of the quantum Fisher information-based figure of merit when the optimization is performed over all pure product states. We eventually show that the difference between the figures of merit obtained using the optimal product and the GHZ states gradually increases with an increasing number of parties.
Theorem 1.
In the estimation of $N$ independent single-body field strengths, given by $h=\{h_{1},h_{2},...,h_{N}\}$, of a many-body arbitrary encoding Hamiltonian, $\widetilde{H}=\hbar\sum_{i=1}^{N}h_{i}H_{i}$, and arbitrary weight matrix, $W$, the quantum Fisher information-based figure of merit, $\text{Tr}(W\widetilde{F}_{Q}^{-1})$, has a lower bound given by
$\text{Tr}(W\widetilde{F}_{Q}^{-1})\geq\frac{\text{Tr}(W^{1/2})^{2}}{\sum_{i=1}^{N}(\lambda_{M}^{i}-\lambda_{m}^{i})^{2}},$ (15)
where $\lambda_{M(m)}^{i}$ denotes the maximum (minimum) eigenvalue of $H_{i}$, $N$ denotes the number of parties, and the minimization is performed over all $N$-party pure input probes.
Proof. Let us suppose that the inverse of the QFIM, $F_{Q}$, exists, and let $W$ be an arbitrary positive semi-definite weight matrix. Then we obtain the following chain of inequalities:
$\text{Tr}(W^{1/2})$ $=$ $\text{Tr}\left(F_{Q}^{1/2}W^{1/2}F_{Q}^{-1/2}\right)$ (16)
$=$ $\text{Tr}\left[\left(F_{Q}^{1/2}\right)^{\dagger}W^{1/2}F_{Q}^{-1/2}\right]$
$\leq$ $\sqrt{\text{Tr}\left[F_{Q}^{1/2}\left(F_{Q}^{1/2}\right)^{\dagger}\right]}$
$\sqrt{\text{Tr}\left[W^{1/2}F_{Q}^{-1/2}\left(W^{1/2}F_{Q}^{-1/2}\right)^{\dagger}\right]}$
$=$ $\sqrt{\text{Tr}(F_{Q})\text{Tr}\left[W^{1/2}F_{Q}^{-1}W^{1/2}\right]}$
$=$ $\sqrt{\text{Tr}(F_{Q})\text{Tr}(WF_{Q}^{-1})}.$
In deriving the above inequalities, we have used the property of positive semi-definite matrices that all real powers of a positive semi-definite matrix are also positive semi-definite. Specifically, in the second step we have used the feature that $F_{Q}^{1/2}$ is Hermitian, since $F_{Q}$, and hence $F_{Q}^{1/2}$, is positive semi-definite. In the third step, we have used the Cauchy-Schwarz inequality, and the fact that $\text{Tr}(W^{1/2})$ is a real quantity since $W^{1/2}\geq 0$. The Cauchy-Schwarz inequality for any two operators, $A$ and $B$, is given by $|\text{Tr}(A^{\dagger}B)|^{2}\leq\text{Tr}(A^{\dagger}A)\text{Tr}(B^{\dagger}B)$. The fourth step follows since both $F_{Q}^{-1/2}$ and $W^{1/2}$ are positive semi-definite, and hence are Hermitian. Therefore from inequation (16), we obtain
$\text{Tr}(WF_{Q}^{-1})\geq\frac{\text{Tr}(W^{1/2})^{2}}{\text{Tr}(F_{Q})}.$ (17)
Our goal is to minimize the quantity on the right hand side of the above inequality, over all pure input probes, for a given but arbitrary $W$. So this essentially reduces to a maximization of $\text{Tr}(F_{Q})$ over pure input states, for an encoding Hamiltonian, $\widetilde{H}=\hbar\sum_{i=1}^{N}h_{i}H_{i}$. Our aim is to simultaneously estimate $h_{i}$, for $i=1$ to $N$. Let the pure input probe that maximizes $\text{Tr}(F_{Q})$ be denoted by |$\ket{\widetilde{{\psi}_{0}}}$. If the corresponding QFIM be denoted by $\widetilde{F}_{Q}$, then
[…]
One can further write Eq. (18) as
$\text{Tr}(\widetilde{F}_{Q})$ $=$ $4\sum_{i=1}^{N}\left[\text{Tr}(H_{i}^{2}\widetilde{\rho_{i}})-\text{Tr}(H_{i}\widetilde{\rho_{i}})^{2}\right]$ (19)
$=$ $4\sum_{i=1}^{N}\Delta^{2}_{\rho_{i}}H_{i},\;\;$
where $\widetilde{\rho_{i}}=\text{Tr}_{\overline{i}}(|\widetilde{{\psi}_{0}}\rangle\langle\widetilde{{\psi}_{0}}|)$, and $\overline{i}$ denotes every other party except the $i^{\text{th}}$ one. Here $\widetilde{\rho_{i}}$ is the reduced density matrix corresponding to the $i^{\text{th}}$ party, which can be, in general, mixed. Since |$\ket{\widetilde{\psi_{0}}}$ is the state that maximizes $\text{Tr}(F_{Q})$, $\text{Tr}(\widetilde{F}_{Q})$ corresponds to the maximum of $\text{Tr}(F_{Q})$. Now, if we consider each term of the summation in Eq. (19), i.e. $\Delta^{2}_{\rho_{i}}H_{i}$, then the maximum of each such term is attained by pure states, and the maximum value is given by $(\lambda_{M}^{i}-\lambda_{m}^{i})^{2}/4$ Textor (1978). The corresponding optimal state that maximizes $\Delta^{2}_{\rho_{i}}H_{i}$ is |$\ket{\chi^{i}}=(\ket{\lambda_{M}^{i}}+\ket{\lambda_{m}^{i}})/\sqrt{2}$, where |$\ket{\lambda_{M(m)}^{i}}$ denotes the eigenvectors corresponding to the maximum (minimum) eigenvalue of $H_{i}$. Therefore the state that maximizes $\text{Tr}(\widetilde{{F}_{Q}})$ is |$\ket{\widetilde{\psi_{0}}}=\ket{\chi^{i}}^{\otimes N}$, which is pure product. From the above analysis, we find that the maximum value of $\text{Tr}(\widetilde{F}_{Q})$ is $\sum_{i=1}^{N}(\lambda_{M}^{i}-\lambda_{m}^{i})^{2}$, when optimized over pure probes. Using inequation (17), this finally leads us to the result that the minimum value of $\text{Tr}(WF_{Q}^{-1})$, when optimized over all pure input probes, is given by $\text{Tr}(W^{1/2})^{2}/\xi$, where $\xi=\sum_{i=1}^{N}(\lambda_{M}^{i}-\lambda_{m}^{i})^{2}$, and $W$ is an arbitrary positive semi-definite matrix, which proves theorem 1. $\blacksquare$
Remark. It is to be noted that the maximum of $\text{Tr}(F_{Q})$ is obtained when |$\ket{\psi_{0}}$ is pure product. However this, in general, is not the state which provides the minimum value of $\text{Tr}(WF_{Q}^{-1})$. For every choice of probe state, the value of $\text{Tr}(WF_{Q}^{-1})$ is always higher than or at most equal to $\text{Tr}(W\widetilde{F}_{Q}^{-1})$. It happens that pure product states maximize the value of $\text{Tr}(\widetilde{F}_{Q})$, and corresponding to these states, the quantity, $\text{Tr}(WF_{Q}^{-1})$ may have a value greater than or equal to $\text{Tr}(W\widetilde{F}_{Q}^{-1})$. This certainly does not imply that $\text{Tr}(WF_{Q}^{-1})$ is minimum for pure product states that maximize $\text{Tr}(F_{Q})$. In fact, in the succeeding subsections, we will show that for a certain choice of the weight matrix, $W$, pure product probes never attain the lower bound of $\text{Tr}(W\widetilde{F}_{Q}^{-1})$.
Corollary 1.
The minimum value of the quantum Fisher information-based figure of merit, given by $\text{Tr}(\mathcal{W}(\mathcal{F}_{Q})^{-1})$, is equal to $1/4$ for the encoding Hamiltonian, $H=\hbar\sum_{i=1}^{N}h_{i}\sigma_{z}^{i}$, and the fixed weight matrix $\mathcal{W}=|\mathcal{W}\rangle\langle\mathcal{W}|$, where |$\ket{\mathcal{W}}=\sum_{i=0}^{N-1}\ket{i}$in the computational basis.
Proof. Considering the encoding Hamiltonian to be $H$, given in Eq. (12), it follows from theorem 1 that the maximum value of $\text{Tr}(\mathcal{F}_{Q})$, optimized over all pure states, is given by $\xi=4N$. Further, if we take $W=\mathcal{W}$, then $\mathcal{W}^{1/2}=\mathcal{W}/\sqrt{N}$, from which it follows that $\text{Tr}(\mathcal{W}^{1/2})=\sqrt{N}$. Therefore the lower bound in (17) for the encoding Hamiltonian, $H$, and weight matrix, $\mathcal{W}$, is given by
$\text{Tr}(\mathcal{W}\mathcal{F}_{Q}^{-1})\geq\frac{1}{4}.$ (20)
So the minimum obtainable value of the quantum Fisher information-based figure of merit, $\text{Tr}(\mathcal{W}\mathcal{F}_{Q}^{-1})$, when optimized over all $N$-party pure input probes is $1/4$ independent of the value of $N$. $\blacksquare$
Corollary 2.
The minimum value of $\text{Tr}(\mathcal{W}(\mathcal{F}_{Q})^{-1})$ for an encoding Hamiltonian, $H=\hbar\sum_{i=1}^{N}h_{i}\sigma_{z}^{i}$, is obtained by the pure probe state which satisfies $\mathcal{F}_{Q}=4\mathcal{W}$, where $\mathcal{F}_{Q}$ is the corresponding QFIM and $\mathcal{W}=|\mathcal{W}\rangle\langle\mathcal{W}|$, where |$\ket{\mathcal{W}}=\sum_{i=0}^{N-1}\ket{i}$ in the computational basis.
Proof. We find the condition under which the equality in (20) is attained. This condition follows from the equality condition of the Cauchy-Schwarz inequality used in the derivation of (16), which is given by $\mathcal{F}_{Q}^{1/2}=\widetilde{K}\mathcal{F}_{Q}^{-1/2}\mathcal{W}^{1/2}$, where $\widetilde{K}$ is a constant independent of $\mathcal{W}$ and the QFIM. This condition essentially implies $\mathcal{F}_{Q}=K\mathcal{W}$, where $K=\widetilde{K}/\sqrt{N}$. Therefore, since $0\leq\text{Tr}(\mathcal{F}_{Q})=\text{Tr}(K\mathcal{W})\leq 4N$, and $\text{Tr}(K\mathcal{W})=NK$, we obtain $0\leq K\leq 4$. Now, using $\mathcal{F}_{Q}=K\mathcal{W}$ in inequality (20), we find that the condition of equality is attained at $K=4$. So, for the encoding Hamiltonian, $H$, the lower bound in (20) is achieved by a pure state for which $\mathcal{F}_{Q}=4\mathcal{W}$. $\blacksquare$
In the subsequent subections, we prove that the particular form of the QFIM, $\mathcal{F}_{Q}=4\mathcal{W}$, is obtainable, under a certain limiting condition, only when the input probe is the GHZ state up to a relative phase, and then find the optimal precision attainable by pure product probes for arbitrary number of parties.
IV.1 Greenberger-Horne-Zeilinger state is the optimal probe
Here we find optimal pure states that provide the best precision in the estimation of independent local fields aligned along the Pauli-$z$ direction, while considering the weight matrix to be $\mathcal{W}$.
Theorem 2.
The limiting value of the quantum Fisher information-based figure of merit, $\lim_{(\alpha,\beta)\to(1,1)}\text{Tr}(\mathcal{W}\mathcal{F}_{Q}^{-1})$, exists which equals to $1/4$, and is attainable only by the GHZ state up to a relative phase, in the eigenbasis of the encoding Hamiltonian $H=\hbar\sum_{i=1}^{N}h_{i}\sigma_{z}^{i}$, where $4\alpha$ and $4\beta$ denotes respectively the diagonal and off-diagonal elements of the QFIM, $\mathcal{F}_{Q}$. The weight matrix considered is $\mathcal{W}=|\mathcal{W}\rangle\langle\mathcal{W}|$, where |$\ket{\mathcal{W}}=\sum_{i=0}^{N-1}\ket{i}$ in the computational basis.
Proof. Let us consider each diagonal and each off-diagonal terms of the QFIM in Eq. (13) and (IV) to be $4\alpha$ and $4\beta$ respectively, i.e. $(\mathcal{F}_{Q})_{ii}=4\alpha$ and $(\mathcal{F}_{Q})_{ij}=4\beta$. From corollary 2, we know that the minimum value of $\text{Tr}(\mathcal{F}_{Q}^{-1}\mathcal{W})$ is obtained when $\mathcal{F}_{Q}$ equals to $4\mathcal{W}$, which would imply $\alpha=\beta=1$. This would in turn imply that $\mathcal{F}_{Q}$ is non-singular. To overcome this problem, let us instead consider the case when $\alpha\neq\beta$, and take the limit $(\alpha,\beta)\to(1,1)$. In such a scenario, we will prove that the quantity,
$\gamma=\lim_{(\alpha,\beta)\to(1,1)}\text{Tr}(\mathcal{F}_{Q}^{-1}\mathcal{W}),$ (21)
exists, and gives a value equal to $1/4$ in the limiting condition.
For a QFIM whose every diagonal term is $4\alpha$ and every off-diagonal term is $4\beta$, it can be shown that
$(\mathcal{F}_{Q}^{-1})_{ii}$ $=$ $\frac{1}{4\alpha-4\beta}\left(1-\frac{\beta}{\alpha-\beta+N\beta}\right),\forall i\;\text{and}$
$(\mathcal{F}_{Q}^{-1})_{ij}$ $=$ $-\frac{\beta}{(4\alpha-4\beta)(\alpha-\beta+N\beta)},\forall i\neq j.$
The derivation is provided in Appendix A. Therefore, it follows that the quantity, $\text{Tr}(\mathcal{F}_{Q}^{-1}\mathcal{W})=N/[4(\beta+\alpha(N-1))]$, where $N$ denotes the total number of parties. So, when we take the limit, $(\alpha,\beta)\to(1,1)$, the quantity, $\text{Tr}(\mathcal{F}_{Q}^{-1}\mathcal{W})$, gives a value equal to $1/4$.
We next prove that the GHZ state, up to a relative phase, in the eigenbasis of the Hamiltonian, suffice to be the only state for which the limit in Eq. (21) is attained. Let us consider the input probe to be |$\ket{\psi_{0}}=\sum_{s}a_{s}\ket{s}$, where $s$ is the binary of 00 to $12^{N}-$ 1. The state, |$\ket{\psi_{0}}$ is normalized, which gives the condition, $\sum_{s}|a_{s}|^{2}=1$. In the limit, $(\alpha,\beta)\to(1,1)$, the quantity, $(\mathcal{F}_{Q})_{ii}=4$, and so we see from (13) that $\lim_{(\alpha,\beta)\to(1,1)}\bra{\psi_{0}}\sigma_{z}^{i}\ket{\psi_{0}}=0$, $\forall i$. Therefore it follows from (IV) that
$\lim_{(\alpha,\beta)\to(1,1)}\bra{\psi_{0}}\sigma_{z}^{i}\otimes\sigma_{z}^{j}\ket{\psi_{0}}=1,\;\;\;\forall i,j.$ (22)
[…]
$\sum_{a_{s}\in\eta}|a_{s}|^{2}-\sum_{a_{s}\in\eta^{\prime}}|a_{s}|^{2}=1.$ (23)
Therefore, from Eq. (23) and the normalization condition, we get $\sum_{a_{s}\in\eta}|a_{s}|^{2}=1$, and $\sum_{a_{s}\in\eta^{\prime}}|a_{s}|^{2}=0$. This suggests that $a_{s}=0$, for all $a_{s}\in\eta^{\prime}$. So only those coefficients, $a_{s}$, are non-zero in which the spins corresponding to the first $(i=1)$ and second $(i=2)$ qubits are either both “up” or both “down”. In a similar manner, we can argue for all values of $i$ and $j$ that, only those coefficients of |$\ket{\psi_{0}}$ are non-zero for which the spins in the $i^{\text{th}}$ and jth positions are either both ‘up” or both “down”. This is possible only if the input probe is of the form, |$\ket{\psi_{0}}=(\alpha_{0}\ket{0}^{\otimes N}+\alpha_{2^{N}-1}\ket{1}^{\otimes N})$. Now we use the condition, $\lim_{(\alpha,\beta)\to(1,1)}\bra{\psi_{0}}\sigma_{z}^{i}\ket{\psi_{0}}=0$, which implies that $|\alpha_{0}|=|\alpha_{2^{N}-1}|$. Finally, the trace of $|\Psi_{0}\rangle\langle\Psi_{0}|$ is equal to one, which gives $|\alpha_{0}|=|\alpha_{2^{N}-1}|=1/\sqrt{2}$. This proves our assertion that only the GHZ state, up to a relative phase, in the eigenbasis of the Hamiltonian, $H$, attains the equality in provided in (20). $\blacksquare$
We therefore prove that the best metrological precision in estimating $N$ independent local fields, while considering the weight matrix, $\mathcal{W}$, is attainable only by the GHZ state, up to a relative phase, in the eigenbasis of the corresponding Hamiltonian. In the succeeding subsection, we further find the minimum value of the quantum Fisher information-based figure of merit, when minimized over all pure product states, as a function of $N$.
IV.2 How much precision can product probes provide?
Our aim is to find the best precision that can be attained in the estimation of multiple independent local fields, each oriented along the Pauli-$z$ direction, while considering the weight matrix, $\mathcal{W}$, if the input probe used is a pure product one.
Theorem 3.
The minimum value of the quantum Fisher information-based figure of merit, in the estimation of $N$ independent local fields corresponding to the encoding Hamiltonian, $H=\hbar\sum_{i=1}^{N}h_{i}\sigma_{z}^{i}$, and weight matrix, $\mathcal{W}$, where |$\ket{\mathcal{W}}=\sum_{i=0}^{N-1}\ket{i}$ in the computational basis, is given by $N/4$, if the optimization is performed over all pure product states.
Proof. The input probe is considered to be a tensor product of $N$ single-qubit pure states, given by
|$\ket{\psi_{0}^{P}}=\prod_{i=1}^{N}\left(\cos\frac{\theta_{i}}{2}\ket{0}+e^{i\phi_{i}}\sin\frac{\theta_{i}}{2}\ket{1}\right),$ (24)
where $P$ in the superscript of |$\ket{\psi_{0}^{P}}$ denotes that the state is product. Our aim is to optimize over this class of states and find the minimum value of the quantum Fisher information-based figure of merit, $\text{Tr}(\mathcal{W}(\mathcal{F}_{Q}^{P})^{-1})$, where $\mathcal{F}_{Q}^{P}$ is the QFIM corresponding to the state, |$\ket{\psi_{0}^{P}}$, and encoding Hamiltonian, $H$. Now, we see that, for input probe |$\ket{\psi_{0}^{P}}$, the off-diagonal terms of $\mathcal{F}_{Q}^{P}$ are all equal to 00, i.e.
$(\mathcal{F}_{Q}^{P})_{ij}$ $=$ $4\Big{[}\bra{\psi_{0}^{P}}\sigma_{z}^{i}\otimes\sigma_{z}^{j}\ket{\psi_{0}^{P}}$ (25)
[…]
for $\forall i,j$ and $i\neq j$. Therefore, $\mathcal{F}_{Q}^{P}$, and hence $({\mathcal{F}_{Q}^{P}})^{-1}$, are diagonal in the computational basis, with the diagonal elements of $({\mathcal{F}_{Q}^{P}})^{-1}$ being given by $(({\mathcal{F}_{Q}^{P}})^{-1})_{ii}=1/[4(1-\text{Tr}(\sigma_{z}^{i}\rho_{i}))^{2}]$, where $\rho_{i}=\text{Tr}_{\bar{i}}|\psi_{0}^{P}\rangle\langle\psi_{0}^{P}|$. Further, since we choose a particular form of the weight matrix, $\mathcal{W}$, we obtain the quantum Fisher information-based figure of merit, $\mathcal{V}$, given by
$\mathcal{V}=\text{Tr}[\mathcal{W}({\mathcal{F}_{Q}^{P}})^{-1}]=\sum_{i=1}^{N}\frac{1}{4[1-\text{Tr}(\sigma_{z}^{i}\rho_{i})^{2}]}.$ (26)
If we evaluate the quantity, $\text{Tr}(\sigma_{z}^{i}\rho_{i})$, corresponding to the state, |$\ket{\psi_{0}^{P}}$, it provides a value equal to $\cos\theta_{i}$. Therefore we obtain
$\mathcal{V}=\sum_{i=1}^{N}\frac{1}{4\sin^{2}\theta_{i}}.$ (27)
Now, we minimize the quantity, $\mathcal{V}$, with respect to |$\ket{\psi_{0}^{P}}$, i.e. minimize the right hand side of Eq. (27) with respect to all $\theta_{i}$. Further, since all $\theta_{i}$ are independent, the minimization can be taken inside the summation, and each $\theta_{i}$ can be minimized independently, which essentially reduces the optimization to
$\min_{\ket{\psi_{0}^{P}}}\mathcal{V}=\sum_{i=1}^{N}\min_{\theta_{i}}\frac{1}{4\sin^{2}\theta_{i}}=\frac{N}{4},$ (28)
where the optimal value of $\theta_{i}=\theta_{i}^{opt}=\pi/2$, $\forall i$. Therefore it is proved that the optimal pure product state gives a minimum value of the quantum Fisher information-based figure of merit equal to $N/4$ for an $N$-party system. The value is obviously larger than that obtained using GHZ state for $N\geq 2$. . $\blacksquare$
Corollary 3.
The optimal pure product probe in the estimation of $N$ independent local fields corresponding to the encoding Hamiltonian, $H=\hbar\sum_{i=1}^{N}h_{i}\sigma_{z}^{i}$, and weight matrix, $\mathcal{W}$, where |$\ket{\mathcal{W}}=\sum_{i=0}^{N-1}\ket{i}$ in the computational basis, is given by $\prod_{i=1}^{N}\left(\ket{0}+\exp(i\phi_{i})\ket{1}\right)/\sqrt{2}$, where $\phi_{i}$ denotes the relative phase of the $i^{\text{th}}$ party.
We therefore conclude from the discussion in the preceding subsections that no pure product probe can attain the lower bound of the quantum Fisher information-based figure of merit provided in inequation (20), and only the GHZ state, up to a relative phase, suffice to be the optimum probe. Moreover, the gap in the minimal value of the quantum Fisher information-based figure of merit between the optimal product probe and the optimal probe increases linearly with the increase in number of probe particles.
V Mixed states do not provide better precision
In this section, we ask the question whether the lower bound of quantum Fisher information-based figure of merit given in theorem (1) is attainable by any mixed state. We find that it is not, i.e. the best precision in the estimation of independent local fields is obtainable only by pure states, and any mixed state provides a precision lower than that.
Theorem 4.
The minimum value of the quantum Fisher information-based figure of merit, $\text{Tr}(WF_{Q}^{-1})$, in the estimation of $N$ single-body independent field strengths of the many-body encoding Hamiltonian, $\widetilde{H}=\hbar\sum_{i=1}^{N}h_{i}H_{i}$, and arbitrary weight matrix, $W$, cannot be attained by any mixed state, i.e. the minimum value obtained using mixed states is larger than that using pure probe states.
Proof. Our aim is to prove that the equality in inequation (15) is achieved only by pure states, while for mixed states, the inequality holds with greater than sign. We begin by considering the the inequality provided in (17), which is true for arbitrary probe states, including mixed states, and arbitrary encodings, corresponding to an arbitrary choice of the weight matrix, $W$. We have to minimize the quantity, $\text{Tr}(W^{1/2})^{2}/\text{Tr}(F_{Q})$, over all arbitrary input states, including mixed ones, corresponding to the encoding Hamiltonian, $\widetilde{H}$. So, we essentially have to find whether there exists any mixed state which maximizes the quantity, $\text{Tr}(F_{Q})$. Let the encoded state after evolving using the Hamiltonian, $\widetilde{H}$, be given by $\rho_{h}=\sum_{i}\eta_{h}^{i}|\eta_{h}^{i}\rangle\langle\eta_{h}^{i}|$. The subscript, $h$, means that the eigenvalues and eigenvectors of the encoded state depend on the set of parameters, $h=\{h_{1},h_{2},...,h_{N}\}$. The diagonal terms of the QFIM in this case are given by
$(F_{Q}^{A})_{ii}=\sum_{\begin{subarray}{c}n,n^{\prime}\\ \eta_{n}+\eta_{n^{\prime}}\neq 0\end{subarray}}4\eta_{n}\frac{|\bra{\eta_{n}}\partial_{i}\rho_{h}\ket{\eta_{n^{\prime}}}|^{2}}{(\eta_{n}+\eta_{n^{\prime}})^{2}},\;\;\forall i,$ (29)
where the $A$ in the superscript in the left hand side denotes that this is the QFIM corresponding to any arbitrary state, which can be, in general mixed. Here $\partial_{i}$ denotes derivative with respect to the $i^{\text{th}}$ field strength, $h_{i}$. So we can write $\partial_{i}\rho_{h}=-i[H_{i},\rho_{h}]$. From this, it follows that
$\bra{\eta_{n}}[H_{i},\rho_{h}]\ket{\eta_{n^{\prime}}}$ $=$ $\bra{\eta_{n}}[H_{i}-\langle H_{i}\rangle_{\rho_{h}},\rho_{h}]\ket{\eta_{n^{\prime}}}$
$=$ $(\eta_{n^{\prime}}-\eta_{n})\bra{\eta_{n}}(H_{i}-\langle H_{i}\rangle_{\rho_{h}})\ket{\eta_{n^{\prime}}}$
Now, using this in Eq. (29), we obtain, $\forall i$,
$(F_{Q}^{A})_{ii}$ $=$ $4\sum_{\begin{subarray}{c}n,n^{\prime}\\ \eta_{n}+\eta_{n^{\prime}}\neq 0\end{subarray}}\eta_{n}\frac{(\eta_{n^{\prime}}-\eta_{n})^{2}}{(\eta_{n^{\prime}}+\eta_{n})^{2}}|\bra{\eta_{n}}\Delta H_{i}\ket{\eta_{n^{\prime}}}|^{2}$
$\leq$ $4\sum_{\begin{subarray}{c}n,n^{\prime}\\ \eta_{n}+\eta_{n^{\prime}}\neq 0\end{subarray}}\eta_{n}r|\bra{\eta_{n}}\Delta H_{i}\ket{\eta_{n^{\prime}}}|^{2}$
$=$ $4r[\langle H_{i}^{2}\rangle_{\rho_{h}}-\langle H_{i}\rangle_{\rho_{h}}^{2}],$
where $\Delta H_{i}=(H_{i}-\langle H_{i}\rangle_{\rho_{h}})$ and $r=(\eta_{M}-\eta_{m})^{2}/(\eta_{M}+\eta_{m})^{2}$. Here $\eta_{M(m)}$ is the maximum (minimum) eigenvalue of $\rho_{h}$. It has been used in deriving (V) that the minimum value of the ratio of eigenvalues is $\min_{n,n^{\prime}}(\eta_{n^{\prime}}/\eta_{n})=\eta_{m}/\eta_{M}$ Beau and del Campo (2017). Now we find the condition at which the inequality (V) becomes an equality, i.e. $(\eta_{n^{\prime}}-\eta_{n})^{2}/(\eta_{n^{\prime}}+\eta_{n})^{2}=r$, $\forall n,n^{\prime}$ such that $\eta_{n}+\eta_{n^{\prime}}\neq 0$. We see that the parameter, $r$, can take a maximum value equal to $1$. This means that corresponding to the maximum value of $r$, the quantity, $(\eta_{n^{\prime}}-\eta_{n})/(\eta_{n^{\prime}}+\eta_{n})=\pm 1$, which in turn implies either $\eta_{n}=0$ or $\eta_{n^{\prime}}=0$, $\forall n,n^{\prime}$ such that $\eta_{n}+\eta_{n^{\prime}}\neq 0$. This suggests that if we consider the pair, $(\eta_{n},\eta_{n^{\prime}})$, one of them is always equal to 00, $\forall n,n^{\prime}$. This can be possible if only one eigenvalue of $\rho_{h}$ is equal to $1$, and all others are 00. It is therefore proved that $(\eta_{n^{\prime}}-\eta_{n})^{2}/(\eta_{n^{\prime}}+\eta_{n})^{2}=r_{max}=1$, only when the encoded state is pure. The quantity, $(F_{Q}^{A})_{ii}$, is maximum at $r=1$, which implies that $\text{Tr}(F_{Q}^{A})$ is also maximum at $r=1$, which is possible only for a pure encoded state. We considered unitary encoding, and thus the input probe is also pure in such a scenario, which proves our claim in theorem 4. $\blacksquare$
If we consider $H_{i}=\sigma_{z}^{i}$ $\forall i$, in inequation (V), then we find that
$\max_{\rho_{h}}\text{Tr}(\mathcal{F}_{Q}^{A})=\max_{\rho_{h}}4r\left[N-\sum_{i}\text{Tr}(\sigma_{z}^{i}\rho_{h})^{2}\right]=4N,\;\;\;$ (31)
where $\mathcal{F}_{Q}^{A}$ is the QFIM corresponding to the encoding Hamiltonian, $H$, and arbitrary input states, including mixed ones. The maximum in Eq. (31) is attained only when the input probe is pure. This is because we proved in theorem 4 that $r=r_{max}=1$, only for a pure input probe. Moreover, we have already seen in the preceding section that among the pure states, the GHZ state up to a relative phase, gives the optimal precision. The GHZ state up to a relative phase, minimizes the quantity, $4\left[N-\sum_{i}\text{Tr}(\sigma_{z}^{i}\rho_{h})^{2}\right]$, and gives a value $4N$. If we now assume the weight matrix to be $W=\mathcal{W}$, it then follows from Eq. (31) that the minimum value of the quantity, $\text{Tr}(\mathcal{W}^{1/2})^{2}/\text{Tr}(\mathcal{F}_{Q}^{A})$, in this case is $N/(4N)=1/4$, which is attainable only when the input probe is pure. Thus we find that the minimum value of the quantum Fisher information-based figure of merit is attainable only by pure probes, and mixed probes provide a precision worse than that using pure ones.
VI What happens if the weight matrix is Identity?
Let us consider the lower bound given in inequation (15). If the weight matrix be $W=I_{N}$, where $I_{N}$ denotes the $N\times N$ Identity matrix, then the inequality will reduce to
$\text{Tr}(\widetilde{F}_{Q}^{-1})\geq\frac{N^{2}}{\sum_{i=1}^{N}(\lambda_{M}^{i}-\lambda_{m}^{i})^{2}},$ (32)
where $\lambda_{M(m)}^{i}$ denotes the maximum (minimum) eigenvalue of the single-body terms, $H_{i}$, of the many-body encoding Hamiltonian, $\widetilde{H}=\hbar\sum_{i=1}^{N}h_{i}H_{i}$. Further, if we consider $H_{i}=\sigma_{z}^{i}$, $\forall i$, then we obtain
$\text{Tr}(\mathcal{F}_{Q}^{-1})\geq\frac{N^{2}}{4N}=\frac{N}{4}.$ (33)
Now, we show that pure product probes can attain the precision equal to $N/4$, which is the minimal value of the quantum Fisher information-based figure of merit, if we consider the weight matrix to be equal to $I_{N}$, i.e. the equality in (33) is achievable by pure product probes. The quantity, $\text{Tr}(\mathcal{F}_{Q}^{-1})$, will have contributions from only diagonal elements of $\mathcal{F}_{Q}$. The diagonal terms of the inverse of $\mathcal{F}_{Q}$ is given by $(({\mathcal{F}_{Q}})^{-1})_{ii}=1/[4(1-\text{Tr}(\sigma_{z}^{i}\rho_{i}))^{2}]$, and therefore $\text{Tr}(\mathcal{F}_{Q}^{-1})$ will just be $\sum_{i=1}^{N}\frac{1}{4[1-\text{Tr}(\sigma_{z}^{i}\rho_{i})^{2}]}$. Then, following an argument similar to as in subsection IV.2, it can be proven that pure product probes of the form $\prod_{i=1}^{N}\left(\ket{0}+\exp(i\phi_{i})\ket{1}\right)/\sqrt{2}$ saturate the inequality in (33). Therefore, we find that, if the weight matrix is considered to be Identity, then the lower bound of the quantum Fisher information-based figure of merit can be attainable by pure product probes. So considering the weight matrix to be Identity, the necessity of genuine multiparty entangled probe is not required for optimal estimation, and even product probes can provide the best precision. However, we have already found that for the choice of weight matrix, $W=\mathcal{W}$, this is not the situation.
VII Conclusion
We analyzed the estimation of multiple independent field strengths of an arbitrary many-body encoding Hamiltonian comprising of one-body terms, and provided a lower bound of a chosen figure of merit, which has been used to gauge the accuracy of estimating the multiple field strengths. Estimation of multiple parameters involve fundamental bounds which are given in terms of matrix inequalities. So to assess the actual precision in estimation, a figure of merit is defined in terms of a weight matrix, which follows certain properties that a suitable multiparameter estimation measure should satisfy. We provided a lower bound of this figure of merit, in terms of an arbitrary weight matrix and an arbitrary but local encoding Hamiltonian comprising of an arbitrary number of parties.
Further, we considered a particular weight matrix, that takes into account both diagonal and off-diagonal elements of the quantum Fisher information matrix. This allowed us to consider not only the errors in the estimation of the different parameters, but also in some sense, the “correlations” or “couplings” between them. Considering the field terms to be along the Pauli-$z$ direction, and for a particular choice of the weight matrix, we showed that the minimum value of the figure of merit is a constant, independent of the number of parties considered. Further, we proved that this lower bound is attainable only by the Greenberger-Horne-Zeilinger state, in the eigenbasis of the Hamiltonian, if the optimization is performed over all pure states. This is intriguing in the sense that even if we estimate multiple independent fields corresponding to a local Hamiltonian, it is necessary that the genuine-multiparty entangled GHZ state is used as the probe to attain the best metrological precision.
The uncertainties of two non-commuting observables endorse quantumness into the system, with a prime example being the Heisenberg principle. Here we find that even if the parameters to be estimated are attached to “highly” commuting terms - comprising of only single-body terms - there is a necessity of a genuine multiparty entangled state as the optimal probe.
We also provided the minimum value of the figure of merit that is obtainable using pure product probes. In fact, the difference between the figure of merit obtained using the optimal product state and the optimal state, gradually increases with increasing number of parties. We also showed that the lower bound that we provided is never attainable by any mixed state, therefore signifying that the GHZ state, up to a relative phase is the only choice as the optimal probe in the estimation of independent field strengths of arbitrary but local encoding Hamiltonians. Finally, we proved that the choice of the identity operator as the weight matrix, and so ignoring the off-diagonal terms in the QFIM, does not necessitate the deployment of genuine multiparty entanglement in input probes for optimal estimation, and the optimal probe can be a pure product. Therefore, the choice of the weight matrix considered, involving both diagonal and off-diagonal elements, plays a crucial role in necessitating genuine multiparty entangled probes as optimal ones.
En route, we also provided an axiomatic basis for the choice of the figure of merit, to illustrate how its minimum value can potentially be useful in quantum estimation of multiple parameters.
Appendix A Inverse of a matrix whose diagonal and off-diagonal elements are $4\alpha$ and $4\beta$ respectively
Let us consider a QFIM whose each diagonal element is $4\alpha$ and each off-diagonal element is $4\beta$. We have to find the matrix elements of the inverse of this matrix. Let the QFIM be
$\mathcal{F}_{Q}=4\alpha\sum_{i=1}^{N}|i\rangle\langle i|+4\beta\sum_{i,j=1\atop i\neq j}^{N}|i\rangle\langle j|$ (34)
Now let us define a matrix, $D$, and a vector, $v$, as
$D$ $=$ $(4\alpha-4\beta)\sum_{i=1}^{N}|i\rangle\langle i|,\;\text{and}$
$v$ $=$ $\sqrt{4\beta}\sum_{i=1}^{N}\ket{i}.$ (35)
Therefore it follows that
$\mathcal{F}_{Q}=D+vv^{T}.$ (36)
The inverse of this matrix can be calculated using the Sherman-Morrison formula. The inverse of $\mathcal{F}_{Q}$ is given by
$(\mathcal{F}_{Q})^{-1}$ $=$ $(D+vv^{T})^{-1}$ (37)
$=$ $D^{-1}-\frac{D^{-1}vv^{T}D^{-1}}{1+v^{T}D^{-1}v},$
where in the second step, we have used the Sherman-Morrison formula. Now, using Eq. (A), we obtain
$D^{-1}vv^{T}D^{-1}$ $=$ $\frac{\beta}{\alpha-\beta}\sum_{i,j=1}^{N}|i\rangle\langle j|,\;\text{and}$
$v^{T}D^{-1}v$ $=$ $\frac{N\beta}{\alpha-\beta}.$ (38)
Now, if we use these equations in Eq (37), it follows that
$(\mathcal{F}_{Q}^{-1})_{ii}$ $=$ $\frac{1}{4\alpha-4\beta}\left(1-\frac{\beta}{\alpha-\beta+N\beta}\right),\forall i\;\text{and}$
$(\mathcal{F}_{Q}^{-1})_{ij}$ $=$ $-\frac{\beta}{(4\alpha-4\beta)(\alpha-\beta+N\beta)},\forall i\neq j.$